# Ultrastrongly-coupled and Directionally-nonreciprocal Magnon-polaritons in Magnetochiral Metamolecules

### v2

@@ Abstract
Discussion – This study reveals that the MCh metamolecule consisting of a YIG cylinder inserted in a Cu chiral structure results in ultra-strongly coupled and nonreciprocal magnon-polariton at room temperature. Table S1 in Supplemental Material compares the temperature, cavity-free, magnon-polariton coupling ratio, and nonreciprocity obtained in this study with those reported in the literature. Ultrastrong coupling in the magnon-polariton is primarily caused by direct interaction between photons in the Cu chiral structure and magnons in the magnetic YIG cylinder via AC magnetic fields in the $x$-direction. The coupling ratio is reduced in the $z$-axis-orientated metamolecule (Fig. S3(c)) because electromagnetically induced AC magnetic fields in the $z$-direction cannot drive magnons under the external DC magnetic field, $\mu_{0}H_{\rm ext}$, in the $z$-direction. This results in indirect coupling between the chiral cavity photon and magnon in the $z$-axis-orientated metamolecule. Therefore, the coupling ratio can be enhanced by finding a suitable metamolecule orientation.
The second origin of the ultrastrong coupling is the photon confinement in the chiral meta-atom. The chiral meta-atom functions as a resonator or sensitizer; therefore, the MCh metamolecule realizes cavity-free magnon-polariton for the propagating microwave in the free space. In addition to controlling individual quantum systems Bayer2017 ; Rajabali2022 , the magnon-polariton in the free space leads to the observation of the polarization-independent optically-moving magnon-polariton. The optically-moving effect in the MCh and ME media exhibits an electromagnetic response equivalent to that of a regular material moving at a relativistic speed Kong ; Asadchy2018 . A free-space measurement would enable us to observe the optically-moving magnon-polariton. The Hermitian, ultrastrongly-coupled, and optically-moving magnon-polariton may give rise to coherent quantum rectification of quantum information.
@@
Utilizing the fundamental mode ($n$ = 1) is a strategy toward achieving a much stronger coupling between magnons and photons –- the deepstrongly-coupled magnon-polariton, i.e., $g/\omega_{\rm a}\geq$ 1, in the metamolecule. This experiment uses the third-order mode ($n$ = 3) of the chiral cavity photon to obtain the highest value of $g/\omega_{\rm a}$ = 0.22; the magnon-photon coupling is thus reduced by $\eta_{3}$ = 1/3, where $\eta_{\rm n}$ is a phenomenologically introduced mode-dependent coefficient responsible for the space-dependence of the microwave magnetic field in the MCh metamolecule. Here, we assume the net magnitude of the AC magnetic field of the chiral photon to be inversely proportional to the number of antinodes (See more detailed in Supplemental Material). In this way, the $n$ = 1 fundamental mode may enhance the magnon-photon coupling by maximizing the electromagnetic induction due to the magnon dynamics. Additionally, according to the theoretical prediction for the metamolecule Chiba2024 , the coupling ratio is evaluated via $g/\omega_{\rm a}\propto 1/\sqrt{\omega_{\rm a}}$. Therefore, further enhancement toward the deepstrong-coupling is possible by decreasing the chiral resonant frequency, $\omega_{\rm a}$, of the fundamental mode by reducing the system size and changing the design of the metamolecule and materials of the constitutive meta-atoms. The deepstrongly-coupled magnon-polariton will lead to mysterious quantum effects, such as the vacuum Bloch-Siegert shift Li2018 , the super-radiant phase transition Kim_arxiv , and quantum squeezing under thermal equilibrium Hayashida2023 .
Conclusion – We demonstrate cavity-free ultrastrong-couplings of the directionally nonreciprocal magnon-polaritons in a single three-dimensional metamolecule at room temperature. The metamolecule exhibits the coupling ratio of $g/\omega$ = 0.22, which is achieved by direct interaction between microwave photons in the Cu chiral structure and magnons in the YIG cylinder. These experimental results are reproduced well via numerical simulations and theoretical consideration. The MCh metamolecule represents a crucial step toward developing the hybrid quantum systems.
@@
This study demonstrates that metamaterials are promising for integrating different quasi-particles. One quasi-particle is embodied by one meta-atom; therefore, the synthesis of metamolecules and metamaterials by combining different meta-atoms leads to quasi-particle integration. The quasi-particle integration is comprehensively classified according to the interaction between polarization $P$ and magnetization $M$. The conventional ME and MCh effects observed in this study are given by the coupling between $P$ and $M$. Contrastingly, the dynamic coupling of time-derivatives of $P$ and $M$ results in a new collective mode corresponding to an MCh electromagnon-polariton, which is still unclear but anticipated to be observed using metamaterials. Furthermore, we focus on magnons in this study; however, other quasi-particles, such as, excitons, phonons, and plasmons, can be integrated in the metamaterials. Thus, this study signifies a pivotal advancement toward expounding quasi-particle “chemistry” using metamaterials.

### v3

@@ Abstract
Discussion – This study reveals that the MCh metamolecule consisting of a YIG cylinder inserted in a Cu chiral structure results in ultrastrongly-coupled and directionally-nonreciprocal MP at room temperature. Table S1 in Supplemental Material SM compares the temperature, MP coupling ratio, and nonreciprocity obtained in this study with those reported in the literature (see also references Zhang2014 ; Bai2015 therein). Ultrastrong coupling in the MP is primarily caused by direct interaction between photons in the Cu chiral structure and magnons in the magnetic YIG cylinder via AC magnetic fields in the $x$-direction. The coupling ratio is reduced in the $z$-axis-orientated metamolecule (Fig. S3(c)) because electromagnetically induced AC magnetic fields in the $z$-direction cannot drive magnons under the external DC magnetic field, $\mu_{0}H_{\rm ext}$, in the $z$-direction. This results in indirect coupling between the chiral resonant photon and magnon in the $z$-axis-orientated metamolecule. Therefore, the coupling ratio can be enhanced by finding a suitable metamolecule orientation.
@@
Utilizing the fundamental mode ($n$ = 1) is another strategy toward achieving a much stronger coupling between magnons and photons –- the deepstrongly-coupled MP, i.e., $g/\omega_{\rm a}\geq$ 1, in the metamolecule. This experiment uses the third-order mode ($n$ = 3) of the chiral resonant photon to obtain the highest value of $g/\omega_{\rm a}$ = 0.22; the magnon-photon coupling is thus reduced by $\eta_{3}$ = 1/3, where $\eta_{\rm n}$ is a phenomenologically introduced mode-dependent coefficient responsible for the space-dependence of the microwave magnetic field in the MCh metamolecule. Here, we assume the net magnitude of the AC magnetic field of the chiral resonant photon to be inversely proportional to the number of antinodes (See more detailed in Supplemental Material SM ). In this way, the $n$ = 1 fundamental mode may enhance the magnon-photon coupling by maximizing the electromagnetic induction due to the magnon dynamics.
Additionally, according to the theoretical prediction for the metamolecule Chiba2024 , the coupling ratio is evaluated via $g/\omega_{\rm a}\propto 1/\sqrt{\omega_{\rm a}}$. Therefore, further enhancement toward the deepstrong-coupling is possible by decreasing the chiral resonant frequency, $\omega_{\rm a}$, of the fundamental mode by increasing the system size and changing the design of the metamolecule and materials of the constitutive meta-atoms. The deepstrongly-coupled MP will lead to mysterious quantum effects, such as the vacuum Bloch-Siegert shift Li2018 and quantum squeezing under thermal equilibrium Hayashida2023 . The Hermitian, ultrastrongly-coupled, and directionally-nonreciprocal MP may give rise to coherent quantum rectification of quantum information.
The second origin of the ultrastrong coupling in this study is the photon confinement in the chiral meta-atom. The chiral meta-atom functions as a resonator or sensitizer; because the resonator is embedded in the metamolecule that supports MPs, the MPs for propagating microwaves in the free-space can be realized using the metamolecule. As shown in Fig. S5 in Supplemental Material SM , our additional numerical simulation using COMSOL verifies the free-space MP with ultrastrong coupling and directional nonreciprocity. The MP in the free-space is a central topic in quantum optics as it allows control over individual quantum systems Bayer2017 ; Rajabali2022 . Moreover, the nonreciprocal MP in the free-space leads to the observation of the polarization-independent optically-moving effects. The optically-moving effect in the MCh and ME media exhibits an electromagnetic response equivalent to that of a regular material moving at a relativistic speed Kong ; Asadchy2018 ; kodama2024 . Therefore, free-space MPs with ultrastrong coupling and directional nonreciprocity enable us to realize synthetic gauge fields, i.e., the Lorentz force acting on the propagating microwaves Sawada05 .
@@
Last but not least, the quasi-particle integration is comprehensively classified according to the interaction between polarization $P$ and magnetization $M$. The conventional ME and MCh effects observed in this study are given by the coupling between $P$ and $M$. Contrastingly, the dynamic coupling of time-derivatives of $P$ and $M$ results in a new collective mode corresponding to an MCh electromagnon-polariton, which is still unclear but anticipated to be observed using metamaterials. Furthermore, we focus on magnons in this study; however, other quasi-particles, such as, excitons, phonons, and plasmons, can be integrated in the metamaterials. Thus, this study signifies a pivotal advancement toward expounding quasi-particle “chemistry” using metamaterials.
Conclusion – We demonstrate ultrastrong couplings of the directionally-nonreciprocal MPs in a single three-dimensional MCh metamolecule at room temperature. The metamolecule exhibits the coupling ratio of $g/\omega$ = 0.22, which is achieved by direct interaction between magnons in the YIG meta-atom and microwave photons confined in the Cu chiral meta-atom as the resonator. These experimental results are reproduced well via numerical simulations and theoretical consideration. The coupling ratio and directional nonreciprocity can be enhanced by changing the metamolecule orientation and the chiral resonance frequency. The MCh metamolecule represents a crucial step toward developing the hybrid quantum systems and quasi-particle “chemistry”using metamaterials. Furthermore, the optically-moving MPs by the metamolecules in the free-space enable us to realize the Lorentz force acting on the propagating microwaves.